OpenStudy (anonymous):

-7(10)-5y=-60 *check answer* /\ -70-5y=-60 -60+-70=-130 y=-130

OpenStudy (anonymous):

mistake made, I think

OpenStudy (anonymous):

-70-5y=-60 -70-5y +70 =-60 +70 try from here.

OpenStudy (anonymous):

oh y=10

OpenStudy (anonymous):

it would be -5y=10, (after adding 70 to both sides) by what would (a single) y be equal to?

OpenStudy (anonymous):

so if its a negative number and it says subtract i turn it into a positive and add? and so i would add 70 to both -5 and 10

OpenStudy (anonymous):

at the step of having -5y=10, you divide by the -5 on both sides, no?

OpenStudy (anonymous):

i went by the steps of how i solved the equation up there

OpenStudy (anonymous):

the answer is y=-2

OpenStudy (anonymous):

yeah but i need to know how though

OpenStudy (anonymous):

because i got -130 so i need to know what i did wrong

OpenStudy (anonymous):

the correct steps are: Given: -70 -5y = -60 adding 70 to both sides: -70 -5y +70 = -60 +70 (to isolate the -5y) simplifying both sides: -5y=10 (the 70's canceled) dividing both sides by -5: -5y /(-5)=10 /(-5) (to isolate the y) simplifying on both sides: y=-2 (the -5's canceled) so that is the answer, y=-2.

OpenStudy (anonymous):

well its telling me i should do 7*10 then so on, well thats how i can figure it out, but yeah your steps make sense in the word format

OpenStudy (anonymous):

oh yeah those steps are right lol

OpenStudy (anonymous):

I am not sure what they tell you, but what I did is the way:)

OpenStudy (anonymous):

if you got any questions about it, please; you are always welcome to ask!

OpenStudy (anonymous):

alright